### v1

# Searching for Accreting Compact Binaries in eROSITA DR1

Abstract
Compact binaries with accreting white dwarfs, neutron stars, or black holes are crucial for understanding accretion physics. In this study, we identify accreting compact binary candidates in the eROSITA DR1 by combining Gaia astrometry and ZTF time-domain photometry. We select accreting compact binary candidates whose X-ray emission is likely accretion-driven, based on their X-ray-to-optical flux ratios and optical colors, and further identify systems with short-period ZTF light curves, thereby ruling out AGNs. The resulting 22 candidates, including two known compact binaries, represent promising candidates for spectroscopic follow-up to confirm their accreting nature and explore their physical properties. Our results demonstrate the effectiveness of combining X-ray-to-optical flux ratio and optical color, and time-domain photometry in selecting accreting compact binaries. The method is scalable and adaptable to future multi-wavelength sky surveys, offering a promising path toward a more complete census of compact binaries in the Galaxy.
show]guwm@xmu.edu.cn
1 Introduction
Compact binaries, pairing a compact object (neutron star, black hole, or white dwarf) with a companion star, are crucial for probing extreme astrophysical processes, including mass accretion and binary evolution (M. A. Abramowicz & P. C. Fragile, 2013; K. A. Postnov & L. R. Yungelson, 2014; D. Belloni & M. R. Schreiber, 2023). The detection of compact binary integrates multiple observational techniques including X-ray observations, which aim to detect those undergoing mass transfer, such as X-ray binaries and cataclysmic variables (CVs) (R. A. Remillard & J. E. McClintock, 2006; B. Warner, 1995; F. Verbunt et al., 1997). X-ray binaries have been the cornerstone of black hole searches, with discoveries relying on X-ray outbursts reaching luminosities of – erg s-1 (J. M. Corral-Santana et al., 2016). However, outburst-based searches lack the sensitivity to identify the numerous Galactic quiescent X-ray binaries (QXBs) with x-ray luminosities of – erg s-1 (J. P. Lasota, 2000). Novel approaches are therefore essential to uncover these faint systems and expand the census of compact binaries.
Studies have employed X-ray observations to search for QXBs in specific regions (D. J. Helfand, 1980; C. O. Heinke et al., 2003). The Chandra Galactic Bulge Survey, led by R. I. Hynes et al. (2011), used optical spectroscopy to characterize X-ray sources, while the Swift Bulge Survey, led by A. Bahramian et al. (2021), utilized X-ray spectroscopy, variability, and optical/infrared counterparts. Despite their sophistication, these surveys cover only —a negligible fraction of the sky even when combined with XMM-Newton/Swift archival data. This highlights the critical need for all-sky surveys to detect faint QXBs.
The all-sky X-ray survey by ROSAT had limited sensitivity and relatively poor positional accuracy (), hindering the identification of optical counterparts to Galactic sources (W. Voges et al., 1999; T. Boller et al., 2016). In contrast, the extended ROentgen Survey with an Imaging Telescope Array (eROSITA) survey is 25 times more sensitive than ROSAT, with significantly improved localization (P. Predehl et al., 2021). The first eROSITA All-Sky Survey (eRASS1), publicly released on 2024 January 31, contains 1 million X-ray point sources (A. Merloni et al., 2024). This enables the systematic search for QXBs across the entire Milky Way.
Unlike X-ray outbursts, the luminosities of QXBs are comparable to CVs and accreting white dwarfs (U. Kolb, 1998; C. J. Hailey et al., 2018). The population of white dwarf systems overwhelmingly outnumbers neutron star and black hole systems, making them statistically dominant in any blind search for quiescent X-ray binaries (Q. Z. Liu et al., 2007; A. F. Pala et al., 2020). The primary focus of this work is the identification of QXBs in eROSITA data, while white dwarf binaries are expected to dominate the sample due to their higher abundance. For all detected compact binaries, the nature of the accretor can be classified through X-ray spectral analysis combined with multi-band photometric properties (D. A. Smith & V. S. Dhillon, 1998).
In eROSITA DR1’s X-ray sources with reliable flux (, 0.2–10 ; 50 counts, S. Saeedi et al.2022), extragalactic sources dominate. The Galactic QXB population in this flux regime suffers significant contamination by active stars/binaries, which can be efficiently filtered using the X-ray-to-optical flux ratio () and color criteria established by A. C. Rodriguez (2024). This methodology has proven highly effective, with A. C. Rodriguez et al. (2025) identifying 127 CVs within 150 through precise eROSITA/Gaia DR3 astrometric matching, while X. Wang & J. Takata (2025) expanded the sample to 412 CV candidates using ZTF light-curve periodicity analysis.
2 Candidates Selection
Our goal is to identify reliable compact binary candidates in eRASS1. To this end, we require that sources in our sample satisfy the following essential criteria:
1.
Accurate positional matching to an optical counterpart.
2.
X-ray emission consistent with accretion processes.
3.
Evidence of short-period photometric variability.
These conditions ensure the selection of candidate characteristic of compact binaries while minimizing contamination from other x-ray source populations.
2.1 Counterparts
Optical counterpart identification in the eROSITA Final Equatorial-Depth Survey suggests that approximately 10% of X-ray sources are Galactic (A. Liu et al., 2022; M. Salvato et al., 2022). While the typical positional uncertainty of eROSITA is 5′′, it remains significantly larger than that of optical catalogs. As a result, the true optical counterpart is not necessarily the closest source when multiple candidates are present near an X-ray detection. However, no systematic study has yet matched optical counterparts to sources in eRASS1.
To prioritize the accuracy of counterpart identification, we adopt the following criterion:
An eRASS1 source must have exactly one optical source within a 10′′ radius.
The 10′′ matching radius is chosen based on eROSITA’s positional accuracy and statistical matching requirements. The average positional uncertainty of eRASS1 sources is typically less than 10′′, and even smaller at high confidence levels (A. Merloni et al., 2024). This radius effectively balances the need to minimize false matches while ensuring the inclusion of true optical counterparts.
The depth of an optical survey determines its ability to detect faint sources, which is critical for identifying the true optical counterparts of X-ray sources. To match the sensitivity of eROSITA, the optical catalog must have a detection limit sufficient to resolve these counterparts. Assuming an X-ray sensitivity of erg s-1cm-2 and typical QXB luminosities of –erg s-1, we estimate that eROSITA can detect such sources out to 3 kpc. At this distance, a companion star with a luminosity of 0.1–1 would have an apparent magnitude of 19–20. Consequently, we adopt Gaia DR3 as the optical catalog for counterpart identification.
Gaia DR3 provides BP/RP photometry for over 1.5 billion sources across the full sky (Gaia Collaboration et al., 2023). We cross-match eRASS1 sources with Gaia DR3 within a 10′′ radius, identifying 367,953 X-ray sources with at least one Gaia match. We then retain only those sources with exactly one Gaia counterpart within this radius. As illustrated in Figure 1, a sky map of a representative source demonstrates the absence of any additional bright Gaia objects within the 10′′ search region (blue circle) centered on the eRASS1 position.
This process yields a list of 367,953 sources matched to optical counterparts. From this, we select 127,083 targets with for subsequent analysis.
2.2 Distinguishing Candidates from Active Stars
Magnetically active stars constitute a dominant class of Galactic X-ray sources (E. D. Feigelson et al., 2002; M. A. Agüeros et al., 2009; V. M. Kaspi & A. M. Beloborodov, 2017). To distinguish X-ray emission due to stellar activity from that produced by accretion, we use Gaia photometry in conjunction with the X-ray-to-optical flux ratio. For the 127,083 sources with Gaia counterparts, we retrieve , , and from Gaia DR3 (phot_g_mean_flux, phot_bp_mean_flux and phot_rp_mean_flux), along with X-ray fluxes in the 0.2–2.3 keV band from eROSITA DR1 (ML_FLUX_1). These values are plotted in the vs. BP–RP diagram shown in Figure 2.
To identify accretion-powered systems, we adopt the empirical boundary proposed by A. C. Rodriguez (2024): . Here, the optical flux is calculated as , and the X-ray flux is used directly from (ML_FLUX_1) without absorption correction. Applying this criterion yields 57,940 sources with flux ratios above the activity threshold.
Given the known discrepancies between Gaia DR3 astrometric solutions and dynamical constraints for some low mass x-ray binaries (LMXBs) (J. Wu et al., 2016; P. Gandhi et al., 2019; P. Charles et al., 2019), we do not impose cuts on astrometric parameters. As a result, the sample includes a number of extragalactic X-ray sources such as AGNs and quasars. However, these contaminants are effectively removed in the subsequent step based on short-period optical variability.
2.3 ZTF Light Curve
Photometric light curves are a powerful diagnostic for detecting periodic variability in quiescent X-ray binaries (QXBs). In such systems, tidal distortion of the companion can induce ellipsoidal modulation, while other features, such as eclipses, reflection effects, or stochastic accretion variability, may complicate or obscure the ellipsoidal modulation. Despite these complexities, time-domain analysis remains essential for identifying periodic behavior indicative of compact binaries.
The Zwicky Transient Facility (ZTF) is a wide-field optical survey covering the northern sky to a depth of 20.5 (r band, 5) (E. C. Bellm et al., 2019). Its 22nd public data release (DR22) includes 58.5 million images and 4.92 billion light curves. The ZTF collection of light curve data is accessed through the Application Programming Interface (API) provided by Infrared Science Archive (IRSA)111https://irsa.ipac.caltech.edu/docs/program_interface/ztf_lightcurve_api.html. For each of the 57,940 X-ray-selected candidates, we query g- and i-band photometry within a 5′′ radius of the Gaia position using the IRSA API. Of the 57,940 sources, 17,221 match ZTF DR22 time-domain photometry, yielding 33,048 light curves in two bands.
We employ the Lomb-Scargle algorithm to search for periodic signals (N. R. Lomb, 1976; J. D. Scargle, 1981), folding the light curves with the obtained periods . The period search window is set to 0.1–4 days, with frequency boundaries defined as minimum_frequency = 0.25 and maximum_frequency = 10. For each candidate, we fold the light curves using the period of maximum power and visually inspect both the power spectrum and phase-folded light curve. This yields 324 candidates exhibiting significant periodic modulation.
These 324 candidates are then visually inspected to eliminate those with periods likely induced by Earth’s rotation (e.g., period very close to 1 day) or improper folding. Ultimately, we select 22 targets. While their light curves may exhibit contamination or additional components, the dominant variability in each case corresponds to the period of the peak power in the Lomb-Scargle spectrum. An example of a periodic light curve is shown in Figure 1.
3 Result
We identify 22 compact binary candidates in eROSITA DR1 through a three-step selection process: (1) cross-matching to a unique Gaia optical counterpart within a 10′′ radius; (2) requiring X-ray-to-optical flux ratios exceeding the empirical threshold for stellar activity; and (3) detecting short-period variability in ZTF time-domain photometry. Among these candidates, two are previously known systems: J041920072545 (hereafter J0419) and J102347003840 (hereafter J1023). J0419 was confirmed via dynamical analysis using data from the Large Sky Area Multi-Object Fiber Spectroscopic Telescope (LAMOST) as a binary comprising an extremely low-mass pre-white dwarf and a compact object with a mass of (X.-Q. Cui et al., 2012; Z.-X. Zhang et al., 2022). J1023 exhibits a bistable accretion behavior, switching between a radio pulsar state and a LMXB phase depending on the mass transfer rate (A. M. Archibald et al., 2009). A summary of all 22 candidates is provided in Table1.
3.1 Archival Spectroscopic Survey Data Matching
We cross-match our candidates with the LAMOST and Sloan Digital Sky Survey (SDSS, R. Ahumada et al.2020). LAMOST provides time-domain spectroscopy of J0419, and reveals double-peaked emission lines indicative of an accretion disk in spectra. Additionally, LAMOST conducted low-resolution spectroscopic observations of J063213253622 (hereafter J0632), J085909053654 (hereafter J0859), and J094558292252 (hereafter J0945), which are classified as CVs in the LAMOST Low Resolution Survey. The SDSS database includes spectroscopic observations of J1023, J0859, and J0945. The spectrum of J1023 exhibits double-peaked emission lines from an accretion disk, while J0859 and J0945 show emission line features characteristic of CVs. No spectroscopic data from either LAMOST or SDSS are available for the remaining 17 candidates.
3.2 SIMBAD Classifications of Candidates
We searched for SIMBAD entries corresponding to our 22 candidates by querying their optical coordinates. SIMBAD has adopted classifications for 15 of 22 candidates, including the well-studied compact binaries J0419 and J1023. Among them, 12 classifications are based on sources from Gaia EDR3, as illustrated in Figure 3. According to Gaia EDR3, J0632, J081210040351, J0859, and J0945 are classified as CVs; J051723115359 and J052316252737 as RR Lyrae stars; and J064513195511, J073223072292, J073644131707, J080038275245, J093757171012 (hereafter J0937), and J094146060933 as eclipsing binaries. Additionally, ZTF classified J0419 and J152611102512 as BY Draconis variables. The remaining 7 candidates currently have no classification or identification recorded in SIMBAD.
3.3 eROSITA X-ray Analysis
For candidates with X-ray fluxes exceeding the threshold , we retrieve their source detection products to fit the X-ray spectra. Using the grppha tool, we group the spectra with a minimum of three channels per bin. We fit the spectra in the 0.3–5 keV range using a simple absorbed power-law model, tbabs*pow, and estimate the unabsorbed flux in the 0.5–10 keV band using the cflux command in XSPEC (version 12.12.0). The tbabs model (J. Wilms et al., 2000) employs photoelectric cross-sections from D. A. Verner et al. (1996). All reported errors are at the 1 confidence level. An example of the fitting on X-ray spectrum is shown in Figure 1.
The limited photon counts hinder robust spectral analysis, resulting in large uncertainties in the fitted parameters and reduced reliability of the tbabs*pow model. This scarcity of data prevents detailed characterization of the X-ray emission properties for most candidates. Future observations with longer exposure times or joint analyses with other X-ray instruments could mitigate these limitations, enabling more precise spectral modeling and deeper insights into the nature of these systems.
In the absence of sufficient photon counts for full spectral fitting of all candidates, we use fluxes from the eROSITA DR1 catalog to construct hardness ratios. The catalog provides fluxes in four energy sub-bands, with ML_FLUX_P1, ML_FLUX_P2, ML_FLUX_P3, and ML_FLUX_P4 respectively representing the 0.2 to 0.5 keV, 0.5 to 1 keV, 1 to 2 keV, and 2 to 5 keV ranges. In subsequent calculations, we denote these fluxes as P1, P2, P3, and P4. We calculate a hardness ratio (HR) for each source to characterize its spectral properties. The hardness ratio is defined as:
We then examine the distribution of HR as a function of orbital period derived from ZTF light curves (see Figure 4). Cataclysmic variables (CVs) are expected to occupy the lower-left region of this diagram, characterized by short orbital periods (8 hours) and relatively soft X-ray spectra. In contrast, neutron star or black hole binaries generally exhibit harder X-ray emission. The distribution of our 22 candidates is broadly consistent with this trend, with the lower envelope showing a positive correlation between HR and orbital period. This suggests that the combination of X-ray hardness and orbital period may provide a useful empirical tool for distinguishing CVs from more compact accretors such as neutron stars and black holes.
3.4 GALEX Ultraviolet Data
We utilize the Galaxy Evolution Explorer (GALEX) Data Release 7 (DR7) to match ultraviolet photometry for our 22 candidate sources. Through cross-matching the GALEX DR7 catalog with our sample within a 5′′ radius, we obtain NUV magnitudes for seven sources, including J0632 and J0859. Additionally, we query the Multi-mission Archive at STScI (MAST) for GALEX tiles upon our candidates and download the associated Coadded Imaging Products. The Merged Catalog of Sources (AIS__mcat.fits) provides NUV magnitudes for other six sources. We note that some faint sources (NUV 23 in the merged catalog) have fluxes approaching the background noise level. for For five faint candidates, we derive flux upper limits of 5 using the original Imaging files (AIS__int.fits). Figure 4 illustrates the distribution of the NUV-to-X-ray flux ratio versus orbital period for our sample. For compact binaries with comparable accretion rates, neutron star or black hole systems are expected to exhibit harder X-ray emission compared to white dwarf systems. The NUV-to-X-ray flux ratios of our candidates span several orders of magnitude. We suggest that sources with higher flux ratios are more likely to be neutron star or black hole X-ray binaries. With additional classified white dwarf and neutron star/black hole systems, we anticipate a clearer separation in this distribution, enabling more precise classification of candidate sources.
Gaia Source | |||||||||
(mas) | () | () | (day) | ||||||
J041920072545 | 1.45 | 52.38 | 14.7 | 0.95 | 0.607 | 2.44 | |||
J051723115359 | 0.48 | 7.88 | 17.01 | 1.05 | 0.612 | – | |||
J052316252737 | 0.45 | 10.87 | 16.53 | 1.19 | 0.688 | 1.45 | |||
J060502243748 | 1.70 | 124.21 | 13.84 | 0.96 | 0.638 | – | |||
J063213253622 | 2.24 | 73.67 | 15.11 | 1.01 | 0.324 | 1.21 | |||
J064513195511 | 1.49 | 23.26 | 16.52 | 1.52 | 0.298 | 0.88 | |||
J071307172149 | 0.45 | 4.04 | 17.98 | 1.75 | 0.495 | – | |||
J072518081255 | 0.77 | 15.91 | 16.22 | 1.05 | 0.508 | – | |||
J073223072292 | 0.33 | 6.21 | 16.45 | 1.11 | 0.375 | – | |||
J073644131707 | 0.61 | 20.82 | 15.35 | 0.95 | 0.343 | – | |||
J074236020118 | 0.63 | 23.50 | 15.06 | 0.16 | 0.298 | 7.55 | |||
J074332061045 | 0.92 | 6.49 | 17.80 | 1.56 | 0.310 | – | |||
J080038275245 | 0.34 | 4.16 | 17.61 | 1.11 | 0.381 | – | |||
J081210040351 | 0.54 | 2.83 | 18.69 | 0.56 | 0.225 | 3.53 | |||
J085909053654 | 2.25 | 19.01 | 18.07 | 0.99 | 0.222 | – | |||
J093757171012 | 2.48 | 132.87 | 13.09 | 0.79 | 0.482 | 1.64 | |||
J094146060933 | 1.54 | 27.81 | 15.98 | 1.43 | 0.410 | 2.24 | |||
J094558292252 | 2.65 | 9.44 | 18.98 | 0.13 | 0.247 | 1.15 | |||
J102347003840 | 0.69 | 9.72 | 16.23 | 0.75 | 0.396 | 2.24 | |||
J134752263155 | 0.45 | 2.61 | 18.18 | 0.99 | 0.370 | – | |||
J152611102512 | 0.59 | 16.93 | 15.16 | 0.79 | 1.101 | 6.27 | |||
J073523123009 | -0.01 | -0.06 | 17.90 | 1.06 | 0.302 | – |
Note. — This table summarizes the properties of 22 candidate compact binaries. Columns include: Gaia source designation; parallax () and its signal-to-noise ratio (); Gaia G-band magnitude () and color (); X-ray flux () and luminosity () from eROSITA; orbital period from ZTF light curves; X-ray spectral index from power-law fitting (if available). A dash (–) indicates insufficient photon counts to perform a reliable fit.
4 Conclusions and Discussion
We combine Gaia DR3 and ZTF time-domain photometry to search for compact binary candidates in the eRASS1 catalog. We focus on those eRASS1 X-ray sources, which match only one Gaia counterparts within a 10′′ radius. Using the optical-to-X-ray flux ratio and Gaia color, we select the X-ray sources whose X-ray emission dominately produced by accretion. We identify 324 sources exhibiting periodic light curves by searching for periodic signals in ZTF time-domain photometry. Finally, we obtain 22 compact binaries candidates after visual inspection.
Among our candidates, 12 sources are classified as RR stars two well-studied compact binaries J0419 and J1023 are included. J0419 is a pre-extremely low mass (pre-ELM) system, and J1023 is a neutron star X-ray binary.
We collect spectroscopic data from the Large Sky Area Multi-Object Fiber Spectroscopic Telescope (LAMOST) and Sloan Digital Sky Survey (SDSS) for five sources: J0419, J1023, J0632, J0859, and J0945. J0419 and J1023 have multiple spectra, with features consistent with their established classifications. J0632, J0859, and J0945 exhibit spectral characteristics of CV and are classified as CVs by LAMOST and/or SDSS. The Gaia EDR3 catalog provides photometric classifications for 12 additional sources, including CVs, RR Lyrae stars, and eclipsing binaries, as shown in Figure 2. We fit eROSITA X-ray spectra, with J0937 exhibiting a relatively hard spectral type due to higher photon counts. Using ultraviolet photometry, we find that J0937 has a high ultraviolet-to-X-ray flux ratio, suggesting it is likely a neutron star or black hole X-ray binary.
Reflecting on the selection process, requiring a unique Gaia optical counterpart within 10′′ is a stringent criterion. In the absence of an official eRASS1 optical counterpart catalog, this approach effectively identifies accurate optical counterparts, though it excludes a significant number of X-ray sources. As an initial effort in our series of studies, we plan to adopt more relaxed matching criteria in future searches. When excluding active stars, we focus on the optical-to-X-ray flux ratio rather than absolute flux values, simplifying calculations such as X-ray flux conversions. However, this simplification shifts the empirical relation in Figure 2, potentially excluding one additional source. We retain the original boundary in our results, noting that sources near this boundary warrant further analysis. Nevertheless, the adopted boundary effectively distinguishes active stars from accretion-powered systems.
In selecting ZTF periodic variability, we retain sources with clear light curves regardless of their variability patterns. Accretion disks are luminous and unstable, leading to complex light curves in accreting systems that may resemble those of eclipsing binaries or pulsating variables. However, the prior exclusion of active stars using the optical-to-X-ray flux ratio significantly reduces the likelihood of contamination by eclipsing binaries or pulsating variables. The short-period variability sources we identify are highly promising X-ray binary candidates. We plan to request spectroscopic observation time with large telescopes to classify the compact objects and measure their masses through follow-up spectroscopy.
We initiated this study with Gaia DR3 as the optical catalog, representing a preliminary effort to mine compact binaries from eROSITA data. However, as demonstrated throughout this work, our selection does not rely on Gaia astrometric parameters, as no parallax, proper motion, or RUWE cuts were applied. Importantly, the only requirement on the optical catalog is the availability of broad-band photometry to compute colors and flux ratios. This makes our approach fully adaptable to other wide-field photometric surveys with similar depth and coverage. In conjunction with X-ray fluxes and time-domain photometry, this strategy offers a scalable pathway to uncover quiescent LMXBs and CVs beyond Gaia’s astrometric reach.
acknowledgments
This work was supported by the National Natural Science Foundation of China under grant 12433007. This work is based on data from eROSITA, the soft X-ray instrument aboard SRG, a joint Russian-German science mission supported by the Russian Space Agency (Roskosmos), in the interests of the Russian Academy of Sciences represented by its Space Research Institute (IKI), and the Deutsches Zentrum für Luft- und Raumfahrt (DLR). The SRG spacecraft was built by Lavochkin Association (NPOL) and its subcontractors, and is operated by NPOL with support from the Max Planck Institute for Extraterrestrial Physics (MPE). The development and construction of the eROSITA X-ray instrument was led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg & ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität Munich also participated in the science preparation for eROSITA. ZTF is a public-private partnership, with equal support from the ZTF Partnership and from the U.S. National Science Foundation through the Mid-Scale Innovations Program (MSIP).